Grade 12th passGeneral Physics

Determine the frictional force of the air on a body og mass 0.25 kg falling with an acceleration of 9.2 m/s^2

Swadesh Rai
9 Years agoGrade 12th pass

2 Answers

Vikas TU
9 Years ago
the frictional force of the air on a body
Weight – F = (.25)(9.2)
  • 0.25 – F =2.3
  •  F = 2.3-0.25
  • F = 1.05 N
Kunal
8 Years ago
Friction force due to air =m (g-a)Friction force due to air =.25 (9.8-9.2)Friction force due to air =0.15 N